OpenStudy (anonymous):

Use the position function to find the velocity and acceleration at time t = 2. S(t) = sqrt(t^2 + 8).

OpenStudy (anonymous):

@zepdrix can u help?

zepdrix (zepdrix):

The position function is given by \(\large s(t)=\sqrt{t^2+8}\). The derivative function of position is velocity. \(\large s'(t)=v(t)\). The second dervative of position, which is the first derivative of velocity, is the acceleration function. \(\large s''(t)=v'(t)=a(t)\).

zepdrix (zepdrix):

So I'll help you along with the first part, I'll bet you can figure out the acceleration based on this. \(\large s'(t)=\dfrac{t}{\sqrt{t^2+8}}\) This is our velocity function, so let's write it like this, \(\large v(t)=\dfrac{t}{\sqrt{t^2+8}}\) They want us to find the velocity at time t=2. \(\large v(\color{royalblue}{2})=\dfrac{\color{royalblue}{2}}{\sqrt{\color{royalblue}{2}^2+8}}\)

zepdrix (zepdrix):

Understand the process? Now take the derivative of your velocity function to find your acceleration function. \(\large v'(t)=a(t)\) and then again, plug t=2 into it to find the acceleration at that time.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i'm taking the derivative of sqrt(t^2 + 8) right? then i plus in 2 ? @zepdrix

zepdrix (zepdrix):

You/I took the derivative of sqrt(t^2 + 8) to get velocity. Now take the derivative of velocity to get acceleration. So you're now taking the derivative of \[\large v(t)=\dfrac{t}{\sqrt{t^2+8}}\]

OpenStudy (anonymous):

okay so after i take the derivative of that i subtitute 2 in it right?

zepdrix (zepdrix):

yes c:

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

@mathsmind here it is

OpenStudy (anonymous):

so i found the derivative of t^2/sqrt(t^2 + 8) and i got 8/t^2 + 8)^3/2 so i substitute the 2 in the t and i got 8/(2^2 + 8)^3/2 which i'm trying to solve now

OpenStudy (anonymous):

so whats the problem?

OpenStudy (anonymous):

i need to solve 8/(2^2 + 8)^3/2 or should i just leave it like that ?

OpenStudy (anonymous):

well either u work this out with ur head or use the calculator to get the final answer...

OpenStudy (anonymous):

I did use the calculation but i'm not getting a right answer

OpenStudy (anonymous):

@zepdrix i'm stuck on this one can u help?

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

\[s(t) = \sqrt{t^2 + 8}\] \[v(t)=\frac{t}{\sqrt{t^2+8}}\] \[a(t) = \frac{1}{(t^2+8)^{\frac{1}{2}}}-\frac{t^2}{(t^2+8)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now solve for v(2), and a(2) you'll get the answer

OpenStudy (anonymous):

so i should subtitute 2 into t?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i got 0

OpenStudy (anonymous):

\[v(2)=\frac{(2)}{\sqrt{(2)^2+8}}\]

OpenStudy (anonymous):

\[v(2)=\frac{2}{2\sqrt{1+2}}=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

\[m.s^{-1}\]

OpenStudy (anonymous):

what does m.s -1 mean?

OpenStudy (anonymous):

it means meter per a second, u have to put units

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so bow i have to solve the problem above?

OpenStudy (anonymous):

\[a(2) = \frac{1}{((2)^2+8)^{\frac{1}{2}}}-\frac{(2)^2}{((2)^2+8)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

okay so i need to solve that

OpenStudy (anonymous):

a(2)=0.192 m/s^2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

but u really need to know how to differentiate those simple equations ok

OpenStudy (anonymous):

okay